# “Infinitely Often” Transcendence of Gamma-Function Derivatives

### v1

@@ Abstract.
Relatively little is known about the arithmetic properties of Gamma-function derivatives evaluated at arbitrary points $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$. In recent work, we showed that elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ are transcendental infinitely often. This result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. Furthermore, for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ we show that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
@@
Gamma-function derivatives; Euler-Mascheroni constant; transcendence.
@@ 1. Introduction
encodes the Gamma-function derivatives $\Gamma^{\left(n\right)}(q)$ as coefficients. Although $\Gamma\left(t\right),\>t\in\mathbb{R}\setminus\mathbb{Z}_{\leq 0}$ is well understood analytically, it appears that relatively little is known unconditionally about the arithmetic nature of $\Gamma^{\left(n\right)}(q)$ at rational points $q$.
For the base case, $q=1$, the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ satisfies the recursion
$\Gamma^{\left(n\right)}\left(1\right)=-\gamma\Gamma^{\left(n-1\right)}\left(1\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(1\right),\quad n\in\mathbb{Z}_{\geq 1},$
where $\gamma=0.577215\ldots$ denotes the Euler-Mascheroni constant and $\Gamma^{\left(0\right)}\left(1\right)=1$. This system may be iterated to yield [^1] See, for example, Choi and Srivastava [1].
@@
where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ can be expressed as a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); and (iii) the arithmetic properties of $\gamma$ and $\zeta\left(n\right),\>n\in\left\{5,7,\ldots\right\}$ remain unsettled.
@@
Although the polynomial forms on the right-hand side of (2) are generally thought to yield transcendental numbers, no single value $\Gamma^{\left(n\right)}\left(1\right)$ is currently known to be transcendental (or even irrational). In fact, the only unconditional transcendence results of which we are aware are:
  • at least one element of the pair $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(2\right)}\left(1\right)\right\}$ is transcendental (which follows immediately from the fact that $\Gamma^{\left(2\right)}\left(1\right)-\left(\Gamma^{\left(1\right)}\left(1\right)\right)^{2}=\zeta\left(2\right)=\pi^{2}/6$); and
  • the values $\Gamma^{\left(n\right)}\left(1\right)$ are transcendental infinitely often (as recently proved by Powers [2]). [^2] Fischler and Rivoal [3] provided a deep conditional transcendence result, showing that the $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental for all $q\in\mathbb{Q}_{>0}$ if the unproved conjecture $\mathbf{E}\cap\mathbf{D}=\overline{\mathbb{Q}}$ is true (where $\mathbf{E}$ and $\mathbf{D}$ are rings of values taken at algebraic points by arithmetic Gevrey series of orders 1-1 and 11, respectively).
@@
The purpose of the present work is to demonstrate that this “infinitely often” property persists (in some form) throughout the entire rational number line (excluding $\mathbb{Z}_{\leq 0}$). Our first result (Theorem 1) shows that for every integer and half-integer $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ are transcendental infinitely often. Then, for rational $q\notin\tfrac{1}{2}\mathbb{Z}$, our second result (Theorem 2) is disjunctive, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
2. Results
2.1. Values of $\boldsymbol{q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}}$
@@
The theorem below generalizes Theorem 1(ii) of Powers [2], which addressed only the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$. The new result, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, is proved by analogous but more complex arguments based on classical identities for $\Gamma\left(t\right)$ along with constraints imposed by the transcendence of $\pi$.
@@
Essentially, repeated application of the Gamma functional equation reduces $\Gamma\left(q\pm t\right)$ to $\Gamma\left(1\pm t\right)$ or $\Gamma\left(\tfrac{1}{2}\pm t\right)$, and Euler’s reflection formula gives closed forms involving $\sin(\pi t)$ or $\cos(\pi t)$, respectively. The resulting Taylor series then possess coefficients that are algebraic linear combinations of distinct powers of $\pi$, each with a non-vanishing highest-degree term. Comparing these coefficients with those obtained by products of the Taylor series in (1) – under the hypothesis that only finitely many $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental – forces infinitely many distinct powers of $\pi$ to lie in a fixed finite-dimensional vector space over the algebraic numbers, thereby yielding a contradiction.
@@
Euler’s reflection formula provides the two identities,
@@
$\Gamma\left(\frac{1}{2}+t\right)\Gamma\left(\frac{1}{2}-t\right)=\frac{\pi}{\cos\left(\pi t\right)},$
@@
with $a_{j,\xi}\in\mathbb{Q}^{\times}$ for all $j\geq 0$ (because the Taylor-series expansions of $\pi t/\sin\left(\pi t\right)$ and $\pi/\cos\left(\pi t\right)$ contain only non-zero rational coefficients given explicitly in terms of Bernoulli and Euler numbers, respectively). Re-expressing $R_{q}\left(t^{2}\right)$ as
$R_{q}\left(t^{2}\right)={\displaystyle\sum_{k=0}^{\infty}}b_{k,q}t^{2k},$
with $b_{k,q}\in\overline{\mathbb{Q}}$ for all $k\geq 0$ and $b_{0,q}=R_{q}\left(0\right)\neq 0$, we then see from (3) and (4) that
@@
denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space. For any $j>n^{*}$, each summand in (6) possesses at least one algebraic factor $\Gamma^{\left(i_{j}\right)}\left(q\right)$ with $i_{j}>n^{*}$ (because $i+\left(2j-i\right)=2j>2n^{*}\Longrightarrow\left(i>n^{*}\right)\vee\left(2j-i>n^{*}\right)$), and therefore lies in $\mathcal{V}$. Thus, $c_{j,q}\in\mathcal{V}$ for all $j>n^{*}$.
@@
At the same time, it follows from (5) that the family $\left\{c_{j,q}\right\}_{j>n^{*}}$ is $\overline{\mathbb{Q}}$-linearly independent. This is because: (i) if there exists a non-trivial $\overline{\mathbb{Q}}$-linear dependence $\lambda_{1}c_{j_{1},q}+\lambda_{2}c_{j_{2},q}+\cdots+\lambda_{M}c_{j_{M},q}$ (with $j_{1}<j_{2}<\cdots<j_{M}$) for arbitrary $M\in\mathbb{Z}_{\geq 2}$, then the term $\pi^{2\left(j_{M}+1-\xi\right)}$ appears only in $c_{j_{M},q}$ (with non-zero algebraic coefficient $b_{0,q}a_{j_{M},\xi}$), forcing $\lambda_{M}=0$ (because otherwise $\pi$ would be algebraic); and (ii) iterating downward similarly forces $\lambda_{M-1}=\lambda_{M-2}=\cdots=\lambda_{1}=0$.
@@
Consequently, for arbitrarily large $M$, the $M$ elements $\left\{c_{n^{*}+1,q},c_{n^{*}+2,q},\ldots,c_{n^{*}+M,q}\right\}$ must be $\overline{\mathbb{Q}}$-linearly independent and lie in $\mathcal{V}$, which is impossible because $\mathcal{V}$ is finite-dimensional. This contradiction confirms that $\Gamma^{\left(n\right)}\left(q\right)$ is transcendental infinitely often. ∎
$2.$ Values of $\boldsymbol{q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}}$

### v2

@@ Abstract.
Relatively little is known about the arithmetic properties of Gamma-function derivatives evaluated at arbitrary points $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$. In recent work, we showed that the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ contains transcendental elements infinitely often. That result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. Moreover, for all such $q$ we derive a lower bound, $\beta\left(N\right)=\max\left\{0,\sqrt{N}-5/2\right\}/N$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$. For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we find the somewhat weaker result that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
@@
Gamma-function derivatives; Euler-Mascheroni constant; transcendence; density bounds.
@@ 1. Introduction
(for sufficiently small $\left|t\right|$) encodes the Gamma-function derivatives $\Gamma^{\left(n\right)}(q)$ as coefficients. Although $\Gamma\left(z\right)$ is well understood analytically for $z\in\mathbb{R}\setminus\mathbb{Z}_{\leq 0}$ (see, e.g., Artin [1]), it appears that relatively little is known unconditionally about the arithmetic nature of $\Gamma^{\left(n\right)}\left(q\right)$ at rational points $q$ (see Rivoal [2, 3]).
For the cases of $q\in\left\{1,\tfrac{1}{2}\right\}$, the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ satisfies the recursions
$\Gamma^{\left(n\right)}\left(1\right)=-\gamma\Gamma^{\left(n-1\right)}\left(1\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(1\right),$
and
$\Gamma^{\left(n\right)}\left(\tfrac{1}{2}\right)=-g\Gamma^{\left(n-1\right)}\left(\tfrac{1}{2}\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\left(2^{k+1}-1\right)\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(\tfrac{1}{2}\right),$
respectively, where $\gamma=0.577215\ldots$ denotes the Euler-Mascheroni constant, $g=\gamma+2\ln\left(2\right)$, and $\Gamma^{\left(0\right)}\left(q\right)\equiv\Gamma\left(q\right)$. These identities can be iterated to yield the systems [^1] See, e.g., Choi and Srivastava [4].
@@
and
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\$ $\vdots\end{array}$ (3)
@@
where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ is a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); [^2] See Apéry [5] and Beukers [6]. and (iii) the arithmetic properties of $\gamma$ and $\zeta\left(n\right),\>n\in\left\{5,7,\ldots\right\}$ remain unsettled. [^3] See Lagarias [7], Rivoal [2, 3], Ball and Rivoal [8], and Fischler, Sprang, and Zudilin [9].
@@
Although the polynomial forms on the right-hand sides of (2) and (3) are generally thought to yield transcendental numbers, no specific value of $\Gamma^{\left(n\right)}\left(1\right)$ or $\Gamma^{\left(n\right)}\left(\tfrac{1}{2}\right)$, for $n\in\mathbb{Z}_{\geq 1}$, is currently known to be transcendental (or even irrational). In fact, we are aware of only two unconditional transcendence results for these systems:
[…]
@@
In the present work, we investigate the “infinitely often” transcendence of $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$. First, Theorem 1 demonstrates that for every positive integer and every half-integer (i.e., for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$), the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements. This analysis leads directly to Theorem 2, which shows that for all such $q$ the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$ is bounded below by
$\beta\left(N\right)=\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N},$
where $\beta\left(N\right)\asymp N^{-1/2}\rightarrow 0$ as $N\rightarrow\infty$. Finally, for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, Theorem 3 provides a weaker (disjunctive) counterpart to Theorem 1, asserting that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
2. Values of $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$
2.1. “Infinitely Often” Transcendental Derivatives
@@ 2. Values of $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$
We begin by showing that for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, transcendental elements appear infinitely often in the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$. This is proved by arguments based on classical identities for $\Gamma\left(t\right)$ along with constraints imposed by the transcendence of $\pi$.
@@
Essentially, repeated application of the Gamma functional equation reduces $\Gamma\left(q\pm t\right)$ to either $\Gamma\left(1\pm t\right)$ or $\Gamma\left(\tfrac{1}{2}\pm t\right)$, and Euler’s reflection formula, in conjunction with the functional equation, gives closed forms involving $\sin(\pi t)$ or $\cos(\pi t)$, respectively. The resulting Taylor series then possess coefficients that are algebraic linear combinations of distinct powers of $\pi$, each with a non-vanishing highest-degree term. Comparing these coefficients with those obtained by products of the Taylor series in (1) – under the hypothesis that only finitely many $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental – forces infinitely many distinct powers of $\pi$ to lie in a fixed finite-dimensional vector space over the algebraic numbers, thereby yielding a contradiction.
@@
Now note that Euler’s reflection formula, in conjunction with the functional equation, provides the identities
@@
$\Gamma\left(\tfrac{1}{2}+t\right)\Gamma\left(\tfrac{1}{2}-t\right)=\frac{\pi}{\cos\left(\pi t\right)},$
@@
with $a_{j,\xi}\in\mathbb{Q}^{\times}$ for all $j\geq 0$ (because the Taylor-series expansions of $\pi t/\sin\left(\pi t\right)$ and $\pi/\cos\left(\pi t\right)$ contain only even powers with non-zero rational coefficients – given explicitly in terms of Bernoulli and Euler numbers, respectively). Re-expressing $R_{q}\left(t^{2}\right)$ as
$R_{q}\left(t^{2}\right)={\displaystyle\sum_{k=0}^{\infty}}b_{k,q}t^{2k},$ (6)
with $b_{k,q}\in\overline{\mathbb{Q}}$ for all $k\geq 0$ and $b_{0,q}=R_{q}\left(0\right)\neq 0$, we then see from (4), (5), and (6) that
@@
denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space. For any $j>n^{*}$, each summand in (8) possesses at least one algebraic factor $\Gamma^{\left(i_{j}\right)}\left(q\right)$ for some $i_{j}>n^{*}$ (because $i+\left(2j-i\right)=2j>2n^{*}\Longrightarrow\left(i>n^{*}\right)\vee\left(2j-i>n^{*}\right)$), and therefore lies in $\mathcal{V}$. Thus, $c_{j,q}\in\mathcal{V}$ for all $j>n^{*}$.
@@
At the same time, it follows from (7) that the family $\left\{c_{j,q}\right\}_{j>n^{*}}$ is $\overline{\mathbb{Q}}$-linearly independent. This is because: (i) if there exists a non-trivial $\overline{\mathbb{Q}}$-linear dependence $\lambda_{1}c_{j_{1},q}+\lambda_{2}c_{j_{2},q}+\cdots+\lambda_{\mu}c_{j_{\mu},q}=0$ (with $j_{1}<j_{2}<\cdots<j_{\mu}$) for arbitrary $\mu\in\mathbb{Z}_{\geq 2}$, then the term $\pi^{2\left(j_{\mu}+1-\xi\right)}$ appears only in $c_{j_{\mu},q}$ (with non-zero algebraic coefficient $b_{0,q}a_{j_{\mu},\xi}$), forcing $\lambda_{\mu}=0$ (because otherwise $\pi$ would be algebraic); and (ii) iterating downward similarly forces $\lambda_{\mu-1}=\lambda_{\mu-2}=\cdots=\lambda_{1}=0$.
@@
Consequently, for arbitrarily large $\mu$, the $\mu$ elements $\left\{c_{n^{*}+1,q},c_{n^{*}+2,q},\ldots,c_{n^{*}+\mu,q}\right\}$ must be $\overline{\mathbb{Q}}$-linearly independent and lie in $\mathcal{V}$, which is impossible because $\mathcal{V}$ is finite-dimensional. This contradiction confirms that $\Gamma^{\left(n\right)}\left(q\right)$ is transcendental infinitely often. ∎
2.2. Lower Bound for Transcendental Densities
Let
$\delta_{q}\left(N\right)=\dfrac{\#\left\{1\leq n\leq N:\Gamma^{\left(n\right)}\left(q\right)\textrm{ is transcendental}\right\}}{N},\quad N\in\mathbb{Z}_{\geq 1}$
denote the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$, with $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ fixed. We now provide a lower bound for this density.
Theorem 2.
For any $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$,
$\delta_{q}\left(N\right)\geq\dfrac{\max\left\{0,\left\lceil\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2\right\rceil\right\}}{N}\geq\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N}=\beta\left(N\right),$
where $\beta\left(N\right)\asymp N^{-1/2}$ as $N\rightarrow\infty$.
Proof.
Fixing $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ and $N\in\mathbb{Z}_{\geq 1}$, consider the set of coefficients $\left\{c_{j,q}\right\}$ for $j\in\left\{0,1,\ldots,\left\lfloor N/2\right\rfloor\right\}$. From (8), we know that each $c_{j,q}$ is a $\overline{\mathbb{Q}}$-linear combination of products $\Gamma^{\left(u\right)}\left(q\right)\Gamma^{\left(v\right)}\left(q\right)$, with $0\leq u,v\leq N$.
Assume that the number of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$ is given by $N\delta_{q}\left(N\right)=\tau$ for some integer $\tau\geq 0$, and let $\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}}$ denote the $\tau^{*}\leq\tau+1$ distinct transcendental numbers within the set $\left\{\Gamma\left(q\right),\Gamma^{\left(1\right)}\left(q\right),\Gamma^{\left(2\right)}\left(q\right),\ldots,\Gamma^{\left(N\right)}\left(q\right)\right\}$, so that every element of this set lies in $\overline{\mathbb{Q}}\cup\left\{\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}}\right\}$. It then follows that every product $\Gamma^{\left(u\right)}\left(q\right)\Gamma^{\left(v\right)}\left(q\right)$ (with $0\leq u,v\leq N$) is contained in the $\overline{\mathbb{Q}}$-vector space
$\mathcal{U}_{\tau^{*}}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{1,\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}},\alpha_{i}\alpha_{i^{\prime}}:1\leq i\leq i^{\prime}\leq\tau^{*}\right\},$
implying $c_{j,q}\in\mathcal{U}_{\tau^{*}}$.
Since the generating set for $\mathcal{U}_{\tau^{*}}$ has cardinality
$1+\tau^{*}+\frac{\tau^{*}\left(\tau^{*}+1\right)}{2}=\frac{\left(\tau^{*}+1\right)\left(\tau^{*}+2\right)}{2}\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2},$
one can see that
$\dim_{\overline{\mathbb{Q}}}\left(\mathcal{U}_{\tau^{*}}\right)\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2}.$
However, we also know (from arguments similar to those employed in the proof of Theorem 1) that the elements $\left\{c_{0,q},c_{1,q},\dots,c_{\left\lfloor N/2\right\rfloor,q}\right\}$ are $\overline{\mathbb{Q}}$-linearly independent, forcing the span of this set to have dimension $\left\lfloor N/2\right\rfloor+1$. Given that this span is contained in $\mathcal{U}_{\tau^{*}}$, it follows that
$\left\lfloor N/2\right\rfloor+1\leq\dim_{\overline{\mathbb{Q}}}\left(\mathcal{U}_{\tau^{*}}\right)\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2}$
$\Longrightarrow 2\left\lfloor N/2\right\rfloor+2\leq\left(\tau+2\right)\left(\tau+3\right)$
$\Longrightarrow\tau\geq\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2.$
Thus,
$N\delta_{q}\left(N\right)\geq\max\left\{0,\left\lceil\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2\right\rceil\right\}\geq\max\left\{0,\sqrt{N}-5/2\right\}$
$\Longrightarrow\delta_{q}\left(N\right)\geq\beta\left(N\right)=\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N},$
where $\beta\left(N\right)\asymp N^{-1/2}$ as $N\rightarrow\infty$. ∎
Although the indicated lower bound ($\beta\left(N\right)$) is unaffected by $q$, this does not imply that $\delta_{q}\left(N\right)$ itself is constant over $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. After all, $\beta\left(N\right)\asymp N^{-1/2}$ converges to 0 in the limit, thus providing a rather weak (and not very informative) lower bound. At the same time, it is worth noting that this asymptotic bound is actually stronger (i.e., larger in magnitude) than the posited asymptotic lower bound of $\widetilde{\beta}\left(N\right)\asymp\ln\left(N\right)/N$ given by Theorem 1(iv) of Powers [10]. [^6] It also may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational odd-$n$ $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$. More importantly, the new bound is unconditional, whereas the earlier one is conditional on an unproved (albeit computationally supported) conjecture.
3. Values of $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$